Use the References to access important values if needed for this question, A student measures the s concentration in a saturated aqueous solution of nickel() sulfide to be 5.33x10-11 M Based on her data, the solubility product constant for nickel(II) sulfide is Submit Answer Retry Entire Group 7 more group attempts remaining
media%2Ff56%2Ff5643304-2171-4646-8523-56
media%2Fb9d%2Fb9d76356-533b-40bd-ac1c-54
media%2F12f%2F12f1f138-9d8e-4570-b1f4-c2
media%2F10c%2F10c5d3fe-77ea-4c4e-8996-a1
Answer #1

NiS     =     Ni2+     +     S2-

Ksp = [Ni2+] [S2-]

Here [S2-] = 5.33 x 10-11 M = [Ni2+]

So,

Ksp = (5.33 x 10-11) (5.33 x 10-11) = 2.84 x 10-21

Cr(OH)3     =     Cr3+     +     3OH-

Ksp = [Cr3+] [OH-]3

Here [Cr3+] = 1.24 x 10-8 M

So, [OH-] = 3 (1.24 x 10-8) = 3.72 x 10-8 M

Ksp = (1.24 x 10-8) (3.72 x 10-8)3
       = (1.24 x 10-8) (5.15 x 10-23)
      = 6.39 x 10-31

The solubility of Cr(OH)3 is 1.27 x 10-6 g/L

Molar mass of Cr(OH)3 = 103 g /mol

So, 103 g of Cr(OH)3 = 1 mol

1 g of Cr(OH)3 = (1/103) mol

(1.27 x 10-6) g of Cr(OH)3 = [ (1.27 x 10-6) / 103 ] mol
                                           = 1.23 x 10-8 mol

So, the solubility of Cr(OH)3 is 1.27 x 10-6 g/L = 1.23 x 10-8 mol/L
                                                                            = 1.23 x 10-8 M

Now,

Cr(OH)3     =     Cr3+     +     3OH-

Ksp = [Cr3+] [OH-]3

Here [Cr3+] = 1.23 x 10-8 M

So, [OH-] = 3 (1.23 x 10-8) = 3.69 x 10-8 M

Ksp = (1.23 x 10-8) (3.69 x 10-8)3
       = (1.24 x 10-8) (5.02 x 10-23)
      = 6.22 x 10-31

The solubility of Ag2CrO4 is 4.23 x 10-2 g/L

Molar mass of Ag2CrO4 = 332 g /mol

So, 332 g of Ag2CrO4 = 1 mol

1 g of Ag2CrO4 = (1/332) mol

(4.23 x 10-2) g of Ag2CrO4 = [ (4.23 x 10-2) / 332 ] mol
                                            = 1.27 x 10-4 mol

So, the solubility of Ag2CrO4 is 4.23 x 10-2 g/L = 1.27 x 10-4 mol/L
                                                                             = 1.27 x 10-4 M

Now,

Ag2CrO4     =     2Ag+     +     CrO4-2

Ksp = [Ag+]2 [CrO4-2]

Here [Ag+] = 2 x (1.27 x 10-4) M = 2.54 x 10-4 M

So, [CrO4-2] = 1.27 x 10-4 M

Ksp = (2.54 x 10-4)2 (1.27 x 10-4)
       = (6.45 x 10-8) (1.27 x 10-4)
      = 8.19 x 10-12